OpenStudy (anonymous):

A curve passes through the point (0, 5) and has the property that the slope of the curve at every point P is twice the y-coordinate of P. What is the equation of the curve? (Use x as the independent variable.)

OpenStudy (amistre64):

f'(x) = 2x i believe

OpenStudy (anonymous):

Amistre, after that can you look at my other question here please, http://openstudy.com/#/updates/4eaf2766e4b0cb917620b8a8

OpenStudy (amistre64):

and it wants the initial condition of f(0)=5

OpenStudy (anonymous):

It's not 2x :(

OpenStudy (anonymous):

It just wants the equation of the curve. Does that mean mx+b form? or does that mean the function?

OpenStudy (amistre64):

the ycoord of P is simply f(x) with any luck ... so i could be reading it a little off

OpenStudy (amistre64):

a line is generally not a curve, but in some text they consider it a "curve"

OpenStudy (amistre64):

f'(x) = 2 f(x) by chance?

OpenStudy (anonymous):

Well the answer format is y(x) = [ box ]

OpenStudy (amistre64):

y(x) = x^2 + C ; when x=0, y = 5 such that C = 5 y(x) = x^2 + 5; and the slope is y'(x) = 2x the slope at x=6 is 12 while the ycoord is 36 so yeah, that aint it

OpenStudy (amistre64):

im thinking its y= e^2x but i gotta dbl chk my hunch

OpenStudy (anonymous):

Not e^2x :(

OpenStudy (amistre64):

y = e^2x + C y' = 2 e^2x fits when x=0, y=5 means the C = 5 \[y = e^{2x} + 5\] is what makes sense to me

OpenStudy (anonymous):

^^not that either.. Hmm, this one is really tricky.

OpenStudy (amistre64):

it is tricky, but that makes it better than algebra :) "slope of the curve at every point P is twice the y-coordinate of P" the derivative at every x is 2f(x) is how i see it f'(x) = 2 f(x) f'(x)/f(x) = 2 ln(f(x)) = 2x+C f(x) = e^(2x+C) this might be getting close

OpenStudy (amistre64):

when x=0; e^C = 5 C = ln(5) f(x) = e^(2x) * e^(ln(5)) f(x) = e^(2x) * 5 f(x) = 5e^(2x) that might be best

OpenStudy (anonymous):

5e^2x was right. Thanks, I think I see how you did it. I'll have to come back to it later and study what you did.